OpenStudy (mojutmnt):

Write the equation of the line that passes through (3, 6) and (0, –3) in slope-intercept form. Wouldn't that be y = 3x + 3? Choices are y = 3x – 3 y = 3x + 6 y = 3x – 6 y = 3x + 3

jimthompson5910 (jim_thompson5910):

does the point (3,6) lie on y = 3x + 3?

jimthompson5910 (jim_thompson5910):

does the point (0,-3) lie on y = 3x + 3?

OpenStudy (mojutmnt):

Doesn't have a graph.

OpenStudy (mojutmnt):

I got y=2x+3

OpenStudy (mojutmnt):

Crap wrong answer!...

jimthompson5910 (jim_thompson5910):

y = 3x + 3 -3 = 3(0) + 3 ... plug in x = 0 and y = -3 -3 = 0 + 3 -3 = 3 Since that's FALSE, (0,-3) does NOT lie on the line y = 3x + 3

OpenStudy (mojutmnt):

Hmm

OpenStudy (mojutmnt):

y = 3x + 6

jimthompson5910 (jim_thompson5910):

make sure both points lie on this line

OpenStudy (mojutmnt):

K one min

OpenStudy (mojutmnt):

y = 3x - 3 is correct.

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (mojutmnt):

Thanks man! ^~^

jimthompson5910 (jim_thompson5910):

np